A given amplifier is capable of presenting three different input resistance of 200Ω , 50KΩ and 1MΩ. If the amplifier's power gain is 1000 and it's voltage gain is 100 , determine which input to use to provide maximum power output and which to use to provide maximum output voltage for a low impedance microphone of 500Ω and 20mV open circuit voltage determine the power and voltage outputs
1
Expert's answer
2020-10-16T10:59:33-0400

Since the open circuit voltage is the total voltage, the total current will be equal to "I_{total}=\frac{U_0}{R_1+R_2}" , where "U_0" is the open circuit voltage, "R_1" is the microphone impedance, and "R_2" is the resistance of the input load, therefore the input load voltage according to the Ohm law will be "U_2=\frac{U_0}{R_1+R_2}\times R_2" , and the input power:

"P_1 = (\frac{U_0} {R_1+R_2}\times R_2) ^2\times \frac {1}{R_2}" "= \frac{U_0^2\times R_2} {R_1^2+2\times R_1\times R_2+R_2^2}" , an output power is equal to "P_2=P_1\times 1000" "=\frac{U_0^2 \times R_2}{R_1^2+2\times R_1\times R_2+R_2^2}\times 1000" , "1000" is a coefficient of increase in power, it is visible that function "P_2" "(R_2)" decreases on "R" therefore its maximum value will be at "R_2=200 \Omega"

"P_2=\frac{0.02^2\times 200}{500^2+2\times 500\times 200+200^2}\times 1000=0.16" mW.

The voltage is equal at the input "U_2=\frac{U_0}{R_1+R_2}\times R_2" , therefore at the output "U_1=U_2\times 100 =\frac {U _ 0} {R _ 1 + R _ 2 }\times R_2\times 100" , the function "U_1 (R_2)" increases by "R" , therefore at "R_2=10^6\Omega" there will be the highest voltage.

"U_1=U_2\times 100=\frac{0.02}{500+10^6}\times 10^6\times 100=2" V.